OpenStudy (anonymous):

A 2 meter-tall astronaut standing on Mars drops her glasses from her nose. a. How long will it take the glasses to reach the ground? Please help

OpenStudy (anonymous):

You should use kinematics equation. \[h = v _{0} * t - \frac{ 1 }{ 2 } * g * t^{2}\] Just remember that g in Mars is different from here on Earth, assuming you're here also. :) http://www.physicsclassroom.com/class/1dkin/u1l6a.cfm

OpenStudy (anonymous):

I am not sure my science teacher is asking me this and I am only in the 6th grade/ Have not learned it yet

OpenStudy (anonymous):

h=v0 * 2meters -1/2 * gravity* 2meters^2

OpenStudy (anonymous):

gravity on mars = 3.7 Is this correct

OpenStudy (anonymous):

Yes. 3.711 m/s². And v0 is zero, since the glasses have no initial velocity.

OpenStudy (anonymous):

0 *2meters -1/2*3.711*2^2

OpenStudy (anonymous):

t is not 2, but the unknown variable. What you do know is h, the distance from the ground, which is 2.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

0 *t -1/2*3.711*2^2

OpenStudy (anonymous):

would my answer by -7.422 or is this positive 7.422 for "t"

OpenStudy (anonymous):

No. You see correctly that the first term is 0*t, but you insist in substituting the second term, which is quadratic in t, with 2, which is wrong. It is the unknown variable, t. So: h=2 and t you want to know.

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

2 = 0*t - 1/2*3.711*t^2

OpenStudy (anonymous):

2 = -1.8555 t^2

OpenStudy (anonymous):

Yes. The only think you have to notice is the sign of h, which we assume to be positive. But the equation would demands you use -h. You may want to think why by yourself. :)

OpenStudy (anonymous):

so my answer is going to be 2=-1.8555t^2 thank you

OpenStudy (anonymous):

-2 = =-1.8555t^2 You should really think about the -2. Bye.